OpenStudy (anonymous):

Medal and Fan! A)Create a quadratic function with two real zeros. B)Solve a system of non-linear equations with a graph. C) Solve a system of non-linear equations with a table. D) Compare linear, quadratic, and exponential growth functions using a table and a graph. Submit the graph and table. E) Identify that exponential functions exceed linear and quadratic functions.

OpenStudy (anonymous):

@Directrix could you please help me?

OpenStudy (anonymous):

@Loser66 could you please help me?

OpenStudy (anonymous):

@Nnesha could you please help me?

OpenStudy (anonymous):

I really wish to learn how to do this and have been trying all day could anyone teach me step by step?

Directrix (directrix):

This appears to be five questions. For the first: A)Create a quadratic function with two real zeros. For the first one, choose two real numbers, say 2 and 5. If they are zeros of the quadratic function, then: (x - 2) and (x - 5) are factors Multiply (x - 2) and (x - 5) and that will be a quadratic with two real zeros. After you crank out y = (x - 2)* (x - 5), post what you get, okay?

OpenStudy (anonymous):

wait x^2-5x-2x-5

OpenStudy (anonymous):

@Directrix is that right?

Directrix (directrix):

>> wait x^2-5x-2x-5 Combine the like terms of -5x and -2x Also, the -5 is not correct. -2 * -5 = 10 y = x^2 - 7x + 10

OpenStudy (anonymous):

Gotcha y=x^2-7x+10 would be for our step A, so how would I step b with the non-linear equations @Directrix ?

OpenStudy (anonymous):

I would break it down into groups right? x^2-7x+10 (-5x+10)(x^2-2x) -5(x-2)+x(x-2) (x-5)(x-2)

OpenStudy (anonymous):

right @Directrix ?

OpenStudy (anonymous):